Find a quadratic equation with integral coefficients having the given roots.
1+√3 , 1-√3
I don't understand.Can you help me?
Thank You!

Answer by edjones(8007)

1+√3 , 1-√3
1+sqrt(3)=x, 1-sqrt(3)=x
(x-1+sqrt(3))(x-1-sqrt(3))=y
y=x^2-2x-2
.
Ed